If the line through (1, 3) and (a, 9) is parallel to 3x-5y=2, what is a?
Please help me.

Answer by Nate(3500)

3x - 5y = 3
-5y = -3x + 3
5y = 3x - 3
y = 3x/5 - 3/5
Parallel to ~> slope = 3/5
P(1,3) and P(a,9)
M = (y2 + y1)/(x2 + x1) = (9 + 3)/(a + 1) = 12/(a + 1)
The slope(M) must equal 3/5.
12/(a + 1) = 3/5
(a + 1)/12 = 5/3
a + 1 = 20
a = 19